OpenStudy (anonymous):

WILL GIVE MEDAL The cost of 8 muffins and 2 quarts of milk is $18. The cost of 3 muffins and 1 quart of milk is $7.50. How much does 1 muffin and 1 quart of milk cost?

OpenStudy (anonymous):

8m + 2q = 18 3m + 1q = 7.50 (-1) 5m + q = 10.5 q = -5m + 10.5 Now let's plug it in to the first (or second, I prefer the second) equation. 3m + (-5m + 10.5) = 7.50 -2m + 10.5 = 7.50 -2m = -3 m = 3/2 <---- That's what the muffin costs, So $1.50. Now we plug back into any equation and solve for q. q = -5m + 10.5 q=-5(3/2) + 10.5 q = 3 1 muffin costs $1.50 1 quart of milk costs $3.00

OpenStudy (anonymous):

@classof2018 If you'd like to, you can go to my website http://hwhelpguy.weebly.com/ and send me a topic on which to make a video about. For example, I could make a detailed video on this topic, Systems of Linear Equations

OpenStudy (anonymous):

@pabloecortez07 ok